OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is perpendicular to the line 2x -3y = 12 and passes through the point (2, 6).

OpenStudy (across):

What's confusing you about this exercise?

OpenStudy (anonymous):

its y=-3/2x+9

OpenStudy (anonymous):

yes^

OpenStudy (anonymous):

to check it plug in 6=-3/2(2)+9

OpenStudy (anonymous):

6=-3+9 6=6

OpenStudy (anonymous):

hope that helped :)